# Insecurity Through Obscurity: Veiled Vulnerabilities in Closed-Source Contracts

### v1

@@ 3. Case Study: the Destroyer Inu Attack
A notable attack on a closed-source smart contract that relies on external inputs for its operation took place on July 1st, ’24, causing a loss of 22 ETH (worth \$51,056 at the time) (0xprincess, 2024). As the attack involves an attacker-created ERC-20 token called “Destroyer Inu”, we use the same name for the attack itself. This case study shows that control-flow obfuscation adopted by the victim contract ends up hiding known vulnerabilities from analysis tools—without stopping real-world attackers—arguably making contracts less secure.
@@
tx.origin phishing attack.   The core issue is a classic vulnerability where tx.origin is compared to a hard-coded trusted address, but hidden behind control flow obfuscation. This vulnerability is known as the tx.origin phishing attack (by Example, 2025), dating back to at least 2016 (PeterBorah, 2016). In such attacks, an attacker tricks the victim into calling a malicious contract, which then calls the victim’s contract; since both calls have the same tx.origin, the malicious smart contract can execute the victim smart contract with the same privilege as the victim herself.
@@
5  if iszero(eq(from, expected)) {revert(0, 0)} // access check
6 // read two bytes from calldata as the jump destination.
7 let data := calldataload(0x84) // suppose it’s 0x0a00...
@@
20 let success := call(gas(), token, 0, ptr, 68, 0, 0)
@@
Figure 1. Pseudo-Yul MEV bot smart contract. The code starting at 0x0a00 contains a vulnerable function call whose input is determined by the calldata, allowing an adversary to transfer any amount of ERC-20 tokens from the MEV bot’s account to their own. Note that jump and jumpdest are not part of Yul syntax but are included for better readability.
[…]
@@
tx.origin phishing attack, obfuscated.   In the Destroyer Inu Attack, the victim contract is obfuscated. We use pseudo-Yul code to illustrate the smart contract’s logic in Fig. 1. First, at the entry point of the MEV bot’s smart contract in lines 3 to 5, it implements an access control mechanism based on the transaction’s origin to prevent unauthorized addresses from interacting with the contract. Then, in lines 7 to 9, the contract proceeds by jumping to a specific place in code, as determined by the calldata supplied by the transaction. This is unlike “standard” smart contracts that match function calls using the so-called function selector defined as the first 4 bytes of the calldata (Contributors, 2025). In this MEV bot smart contract specifically, when the two calldata bytes at positions 0x85 and 0x86 are 0x0a00, the program jumps to the code segment starting at jumpdest 0x0a00. The functionality of this segment is to manage the transfer of a specific amount of an ERC-20 token to a target address. Therefore, it loads the address of the ERC-20 token, the recipient address, and the amount from the calldata (lines 12 to 14). Then, in lines 15 to 21, it constructs the input based on these parameters and invokes the ERC-20 transfer.

### v3

@@ Appendix B Case Study: the Destroyer Inu Attack
A notable attack on a closed-source smart contract that relies on external inputs for its operation took place on July 1st, ’24, causing a loss of 22 ETH (worth \$51,056 at the time) [1]. As the attack involves an attacker-created ERC-20 token called “Destroyer Inu”, we use the same name for the attack itself. This case study shows that control-flow obfuscation adopted by the victim contract ends up hiding vulnerabilities from analysis tools — without stopping real-world attackers — arguably making contracts less secure.
@@
tx.origin phishing attack.  The first issue is a classic vulnerability where tx.origin is compared to a hard-coded trusted address. This vulnerability is known as the tx.origin phishing attack [8], dating back to at least 2016 [53]. In such attacks, an attacker tricks the victim into calling a malicious contract, which then calls the victim’s contract; since both calls have the same tx.origin, the malicious smart contract can execute the victim smart contract with the same privileges as the victim herself.
A more severe issue that follows the incorrect access-control check is improper asset management. However, existing tools miss the full exploit because of control-flow obfuscation. Existing tools such as Mythril flag it only as a “low-risk” issue. We will discuss the full vulnerability next.
@@
5 // access check
6 if iszero(eq(from, expected)) {revert(0, 0)}
7 // read two bytes from calldata
8 // as the jump destination.
9 // suppose it’s 0x0a00...
10 let data := calldataload(0x84)
@@
23 let success := call(
24 gas(), token, 0, ptr, 68, 0, 0
25 )
@@
Figure 7: Pseudo-Yul MEV bot smart contract. The code starting at 0x0a00 contains a vulnerable function call whose input is determined by the calldata, allowing an adversary to transfer any amount of ERC-20 tokens from the MEV bot’s account to their own. Note that jump and jumpdest are not part of Yul syntax but are included for better readability.
@@
tx.origin phishing attack, obfuscated.  In the Destroyer Inu Attack, the victim contract is obfuscated. We use pseudo-Yul code to illustrate the smart contract’s logic in Fig. 7. First, at the entry point of the MEV bot’s smart contract in lines 3 to 5, it implements an access control mechanism based on the transaction’s origin to prevent unauthorized addresses from interacting with the contract. Then, in lines 7 to 9, the contract proceeds by jumping to a specific place in code, as determined by the calldata supplied by the transaction. This is unlike “standard” smart contracts that match function calls using the so-called function selector defined as the first 4 bytes of the calldata [13]. In this MEV bot smart contract specifically, when the two calldata bytes at positions 0x85 and 0x86 are 0x0a00, the program jumps to the code segment starting at jumpdest 0x0a00. The functionality of this segment is to manage the transfer of a specific amount of an ERC-20 token to a target address. Therefore, it loads the address of the ERC-20 token, the recipient address, and the amount from the calldata (lines 12 to 14). Then, in lines 15 to 21, it constructs the input based on these parameters and invokes the ERC-20 transfer.